pH and Percent Ionization of a Weak Base 20 of 22 > A Review Constants Periodic Table The degree to which a weak base dissociCalculate the percent ionization of HA in a 0.10 M solution. Express your answer as a percent using two significant figures. View Available Hint(s) nothing % Part B Calculate the percent ionization of HA in a 0.010 M solution. Express your answer as a percent using two significant figures. View Available Hint(s)
Answer #1

Part B:

pH = 11.28

pH + pOH = 14.00

pOH = 14.00 -pH

pOH = 14.00 - 11.28

pOH = 2.72

pOH = -log[OH-]

[OH-] = 10^-pOH

[OH-] = 10^-2.72

[OH-] = 0.001905 M

Percent ionization =( [OH-] equilibrium / [NH3]intial ) × 100@%

=( 0.001905 / 0.205) × 100%

= 0.929 %

Hence the percent ionization is 0.929 %

for acid,

percent ionization = ([H3O+] equilibrium / [HA] initial ) × 100%

Actual here Ka of weak acid HA is not given.so unable to calculate the [H3O+] equilibrium